OpenStudy (watchmath):

A fun classic problem Compute \[\int_0^{\pi/2}\frac{\sin^{2011}x}{\sin^{2011}x+\cos^{2011}x}\] If you have seen this problem before, please let others try first :).

OpenStudy (watchmath):

I forgot the dx at the end :)

OpenStudy (anonymous):

Oh I think I have solved a similar problem to this.

OpenStudy (anonymous):

The approach I used to solve that problem was: \[\int\limits_{0}^{a}{f(x) \over f(x)+f(a-x)}dx=\int\limits_{0}^{a}{f(a-x) \over f(x)+f(a-x)}dx\]

OpenStudy (anonymous):

Should I proceed or let others try?!

OpenStudy (watchmath):

That is already a solution I guess :D

OpenStudy (anonymous):

I guess.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

i havea guess. let me see if can prove it

OpenStudy (anonymous):

@Michael: Thanks! :D